The max magnetic field strength of an electromagnetic field is 5 x 10^-6 T. Calculate the maximum electric field strength if the wave is traveling in a medium in which the speed of the wave is 0.75c


1
Expert's answer
2022-01-27T15:17:28-0500

"E=v_0B=\frac{vB}{\sqrt{1-\frac{v^2}{c^2}}}=1.7~\frac{kN}C."